OpenStudy (anonymous):

What is the solution set of the given equation?4y - 6y + 9y = -2

OpenStudy (texaschic101):

first combine all like terms on the right side. Now what do you get ?

OpenStudy (texaschic101):

oops...left side

OpenStudy (anonymous):

could you just tell me the answer?

OpenStudy (texaschic101):

4y - 6y + 9y = -2 13y - 6y = -2 7y = -2 y = -2/7 check.. 4(-2/7) - 6(-2/7) + 9(-2/7) = -2 -8/7 + 12/7 - 18/7 = -2 -26/7 + 12/7 = -2 - 14/7 = -2 -2 = -2 (correct) y = -2/7

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)